# Recurrent Aligned Network for Generalized Pedestrian Trajectory Prediction

### v1

@@ 4.4 Ablation Study
[Uncaptioned image] Figure 4: Ablation study of different alignment approaches on ETH-UCY dataset. The lower the better. [Uncaptioned image] Figure 5: Ablation study of using different loss coefficients on ETH-UCY dataset. X-axis is $\lambda_{1}/\lambda_{2}$. Y-axis is ADE or FDE.
5 Qualitative Analysis
Feature Space Visualization. As illustrated in Figure 6, we visualize the feature spaces under different domains using t-SNE. “w/ RA” and “w/o RA” denote with and without the RA module, respectively. In each figure, we randomly select a test batch data for each domain. Each point in the figure represents the feature of one trajectory. Points in different colours denotes features in different domains. Among these three sets of figures, there is a notably higher feature overlap when using the RA module, demonstrating that our alignment strategy can learn generalized trajectory features. Notably, due to dense interaction, NBA has a more extensive domain gap with ETH-UCY and SDD. The results show that our approach makes feature alignment more obvious in the face of more significant domain gaps, such as ETH-UCY & NBA and SDD & NBA. Predicted trajectory Visualization. As shown in Figure 7, We compare the visualized trajectories with the best performance methods, SocialVAE [40] and Graph-TERN [1], on ETH-UCY dataset. The model is trained on the source domains SDD and NBA. The results in five subsets show that our method surpasses the trajectory prediction accuracy of both SocialVAE and Graph-TERN. This enhanced accuracy is particularly evident in the finer details of the predicted trajectories. This enhancement in prediction accuracy after domain change can be attributed to the effective feature alignment employed in our approach.
@@ 5 Qualitative Analysis
Figure 6: Visualization of feature spaces of different domains using t-SNE. “A & B” on the top row denotes that the source domains are A and B during training. The blue and red dots denote feature spaces of corresponding source domain A and B, respectively. Deep colour regions indicate points overlap.
Figure 7: Visualization of predicted trajectories compared with SocialVAE [40] and Graph-TERN [1] on ETH-UCY dataset.
6 Conclusion
@@
In this paper, we propose and study a new task named generalized pedestrian trajectory prediction. This task aims to generalize the model to unseen domains without accessing the target domain trajectories in the training process. Furthermore, we propose the Recurrent Aligned Network (RAN), a novel alignment framework tailed for generalized pedestrian trajectory prediction. In RAN, we devise a recurrent alignment module with a recurrent alignment loss to effectively align different trajectory domains at both the time-state and time-sequence levels. In addition, we devise a pre-aligned representation module to consider human interactions during the domain alignment process. The experimental results unequivocally highlight the superior generalization capabilities of our method. To the best of our knowledge, our study is the first to explore the generalized setting for pedestrian trajectory prediction, offering significant insights for real-world autonomous driving systems. Limitation and Future Work. One major limitation of our method is that we need two source domains in the training phase due to the characteristics of the domain alignment approach. In the future, we plan to tackle this problem by investigating a single-domain generalization solution for generalized pedestrian trajectory prediction.

### v2

@@ IV-C Quantitative Analysis
Figure 4: Ablation study of different alignment approaches on ETH-UCY dataset. The model is trained on SDD and NBA. The lower the better.
IV-D Ablation Study
In this section, we first study the contribution of the important modules of RAN. Then we evaluate the impact of different alignment approaches. Subsequently, we study the impact of different alignment units and coefficients of our loss function. Finally, we evaluate the efficiency of our method.
Contribution of Each Component. We conduct experiments to evaluate the major components of our RAN on the ETH-UCY dataset, as shown in Table VI. The recurrent alignment module and the pre-aligned representation module are the two major contributions of our proposed framework. PAR and RA denote the pre-aligned representation module and the recurrent alignment module, respectively. To isolate the effect of each component and determine its significance to the overall performance of the model, removing PAR denotes that we only embed the target trajectory for recurrent alignment at each timestamp. Furthermore, we remove RA to evaluate our recurrent aligned strategy.
Figure 5: Ablation study of using different loss coefficients on ETH-UCY dataset. The model is trained on SDD and NBA. X-axis is $\lambda_{1}/\lambda_{2}$. Y-axis is ADE or FDE. The lower the better.
The experimental results show that removing each component results in performance degradation. Specifically, removing PAR results in a performance decrease from 0.23/0.42 to 0.26/0.47 on average ADE/FDE metrics. Removing RA results in a performance decrease from 0.23/0.42 to 0.27/0.50 on average ADE/FDE metrics. Removing both PAR and RA results in a performance decrease from 0.23/0.42 to 0.30/0.55 on average ADE/FDE. These results underscore that both the recurrent alignment module and pre-aligned representation module are effective in improving the generalization capability of our method.
@@ IV-D Ablation Study
Figure 6: Visualization of feature spaces of different domains using t-SNE. “A & B” on the top row denotes that the source domains are A and B during training. The blue and red dots denote feature spaces of the corresponding source domains A and B, respectively. To differentiate point overlap, we set the transparency value of points to 0.3. Deep colour regions indicate points overlap.
Figure 7: Visualization of predicted trajectories compared with SocialVAE [1] and Graph-TERN [63] on ETH-UCY dataset. The model is trained on SDD and NBA. Given the observed trajectories, we illustrate the ground truth paths and predicted trajectories by ours, SocialVAE and Graph-TERN in four different scenes. We see that our results are much closer to the ground truth compared with the other two approaches.
Analysis of Different Source Domain. As shown in Table VII, we conduct experiments to use different source domains in training to validate the effectiveness of our method. The experimental results demonstrate that our method achieves the best performance under all three settings.
The reason for our superior performance in “Case 1” is that, even within the same trajectory dataset, “out-of-distribution” data can exist. These out-of-distribution data are caused by the long-tail distribution phenomenon in trajectory data. Our approach can mitigate the domain shift among out-of-distribution data, leading to improved performance.
The reason for our superior performance in “Case 3” is that we address the domain shift problem by employing a recurrent aligned method specifically tailored to trajectory data. This approach aligns the feature spaces of different domains, enabling the encoder to extract common features, which enhances the model’s generalization capability.
Comparing “Case 1” with “Case 2,” the experimental results show that a small expansion of the training data (from ETH-UCY to ETH-UCY & SDD) leads to a slight performance improvement. This is because both “Case 1” and “Case 2” include the ETH-UCY dataset in their training sets, allowing the model to learn from data in the same domain as the test set. “Case 2” further expands this by adding additional data, serving as data expansion, which accounts for the slight performance enhancement.
Comparing “Case 1” with “Case 3,” the experimental results show that “Case 1” performs better. This is because, although “Case 3” has a larger amount of data, its training set does not include data from the same domain as the ETH-UCY test set, meaning it has not encountered data from the same domain during training. In contrast, both the training and test sets in “Case 1” come from the same domain. Thus, “Case 1” shows better performance than “Case 3”.
Additionally, the performance drop of our method before and after the domain shift is the smallest compared to other methods, further demonstrating the effectiveness of our approach in addressing the domain generalization problem.
TABLE IX: Ablation study about numbers of parameters and inference time for 100 trajectories on NBA dataset.
Inference time # of Parameters
SocialVAE [1] 13.6ms 2.144M
RAN (Ours) 1.9ms 2.222M
Analysis of Different Alignment Approaches. As shown in Figure 4, we replace our recurrent alignment strategy in RAN with time-state and time-sequence alignment on the ETH-UCY dataset. “State” denotes the time-state alignment strategy, implemented by aligning the feature space in each step of the attention layers. “Sequence” denotes the time-sequence alignment strategy, implemented by only aligning the feature space at the last step in the RNN layers. “Recurrent” denotes our recurrent alignment strategy, which considers both time-state and time-sequence alignment.
Experiment results show that our recurrent alignment strategy achieves the best performance compared to the time-state and time-sequence alignment strategies. Specifically, compared to the time-state alignment strategy, we improve the average ADE/FDE from 0.26/0.46 to 0.23/0.42. Compared with the time-sequence alignment strategy, we improve the average ADE/FDE from 0.25/0.44 to 0.23/0.42. This indicates the superiority of our method in domain alignment for generalized pedestrian trajectory prediction. Furthermore, the performance of the time-sequence strategy exceeds that of the time-state alignment strategy, demonstrating the importance of extracting temporal information for pedestrian trajectory prediction tasks.
Analysis of Recurrent Alignment Loss. As shown in Table VIII, we conduct experiments to evaluate the impact of using different alignment losses as the alignment unit in $\mathcal{L}_{rec}$ on ADE/FDE metrics on ETH-UCY dataset. Specifically, MMD, CORAL, KLD, JS, and COS denote using MMD loss, CORAL loss, Kullback-Leibler divergence loss, Jensen–Shannon divergence loss, and cosine similarity as the alignment unit in $\mathcal{L}_{rec}$, respectively.
The evaluation results show that the performance of using the $L_{2}$ loss exceeds that of using other loss functions. It suggests that the $L_{2}$ distance is a more suitable domain distance measurement metric for pedestrian trajectory prediction tasks. As mentioned in Section III, we use the $L_{2}$ distance as the alignment unit in $\mathcal{L}_{rec}$ in this work. One possible explanation is that high-dimensional feature representations may still preserve spatial-level information, i.e., the spatial distance, in pedestrian trajectory prediction tasks.
Analysis of Loss Coefficient. The overall loss function of our RAN consists of two components, recurrent alignment loss $\mathcal{L}_{rec}$ and prediction loss $\mathcal{L}_{pre}$. Since improving generalization can lead to decreased accuracy in the source domain, the weight coefficients for $\mathcal{L}_{rec}$ and $\mathcal{L}_{pre}$ represent the balance between generalization and accuracy in the model, which is important for model training.
As shown in Figure 5, we evaluate different loss coefficients to validate the appropriate proportion of the alignment module in our framework on the ETH-UCY dataset. The proportion of the alignment module is represented by the figure’s loss coefficient ratio $C=\lambda_{1}/\lambda_{2}$. The evaluation results indicate that an excessive proportion of alignment modules sacrifices the model’s prediction accuracy, resulting in a performance decline. Conversely, a deficient proportion of alignment modules diminishes the model’s generalization capacity, leading to a performance decrease. Based on the experimental results, we choose $C=1$ in this work, which achieves the best performance and represents a balance between generalization and precision for the pedestrian trajectory prediction task. In practical applications, if the target scenario is highly similar to the source domain, we choose a smaller $\lambda_{1}$ and a larger $\lambda_{2}$. Conversely, if the target scenario differs significantly from the source domain, we select a larger $\lambda_{1}$ and a smaller $\lambda_{2}$.
Analysis of Efficiency. We conduct experiments to compare the inference time and the number of parameters with SocialVAE [1] on NBA dataset. Since the inference time for each trajectory is very short, our experimental results provide the inference time for 100 pedestrian trajectories.
The experimental results show that while our model has a comparable number of parameters to SocialVAE, our inference time is significantly shorter. The reason for the smaller number of parameters in our model is that, although we require two sets of stepwise attention layers and RNN layers for the two different source domains during training, they share the same parameters. There are two reasons for our faster inference time. First, during the inference phase, we only need a single set of stepwise attention layers and RNN layers for the target domain. Second, our network is more lightweight. Specifically, we use the mixture of expert decoders composed of multiple MLPs to generate multimodal trajectories, rather than using the variational auto-encoder as in previous methods. These experiment results demonstrate that our method is not only effective but also simple and efficient.
V Qualitative Analysis
Feature Space Visualization. As illustrated in Figure 6, we visualize feature spaces in different domains using t-SNE. “w/ RA” and “w/o RA” denote with and without the RA module, respectively. In each figure, we randomly select a test batch data for each domain. Each point in the figure represents the feature of one trajectory. Points in different colours denote features in different domains. Among these three sets of figures, there is a notably higher feature overlap when using the RA module, demonstrating that our alignment strategy can learn generalized trajectory features across different domains. Notably, due to dense interaction, NBA has a more extensive domain gap with ETH-UCY and SDD. The results show that our approach makes feature alignment more obvious in the face of more significant domain gaps, such as ETH-UCY & NBA and SDD & NBA. This also demonstrates the significant generalization capability of our method.
Predicted Trajectory Visualization. As shown in Figure 7, we compare the visualized trajectories in four different scenes with the best performance methods, SocialVAE [1] and Graph-TERN [63], on ETH-UCY dataset. The model is trained in the source domains SDD and NBA. The results in four different scenes show that our method surpasses the trajectory prediction accuracy of both SocialVAE and Graph-TERN. After transferring from the source domain to the target domain, the predictive performance of both SocialVAE and Graph-TERN decreases significantly. Although they can still predict the general trend of future movements, their accuracy is insufficient. In contrast, our method can still accurately predict future trajectories even after domain changes. This enhanced accuracy is particularly evident in the finer details of the predicted trajectories. This enhancement in prediction accuracy after domain change can be attributed to the effective feature alignment employed in our approach.
VI Conclusion
@@
In this paper, we propose and study a new task named generalized pedestrian trajectory prediction. This task aims to generalize the model to unseen domains without accessing the target domain trajectories in the training process. Furthermore, we propose the Recurrent Aligned Network (RAN), a novel alignment framework tailed for generalized pedestrian trajectory prediction. In RAN, we devise a recurrent alignment module with a recurrent alignment loss to effectively align different trajectory domains at both the time-state and time-sequence levels. In addition, we devise a pre-aligned representation module to consider human interactions during the domain alignment process. The experimental results unequivocally highlight the superior generalization capabilities of our method. To the best of our knowledge, our study is the first to explore the generalized setting for pedestrian trajectory prediction, offering significant insights for real-world autonomous driving systems.
Limitation and Future Work. One major limitation of our method is that we need two source domains in the training phase due to the characteristics of the domain alignment approach. In the future, we plan to tackle this problem by investigating a single-domain generalization solution for generalized pedestrian trajectory prediction.